# Maximizing the Maximum Degree in Ordered Nearest Neighbor Graphs

### v2

@@ 1 Introduction
For a given point set $P$ in the plane, in $d$-dimensional Euclidean space, or, even more generally, in an arbitrary metric space, its Nearest Neighbor Graph is a directed graph based on minimum distances. More precisely, these points are the vertices of the Nearest Neighbor Graph, and each vertex has precisely one outgoing edge towards its closest neighbor. To make this notion well-defined, we assume that our point set is in general position, which here means that no point triple determines an isosceles triangle.
These graphs play an important role in modern Computational Geometry due to their relevance in the context of computing geometric shortest paths [15, 16]. Further applications of Nearest Neighbor Graphs for computing spanners, well-separated pairs, and approximate minimum spanning trees in $\mathbb{R}^{d}$ can be found in the survey [20], books [9, 18], or monograph [17], see also [13]. A systematic study of the basic combinatorial properties of Nearest Neighbor Graphs dates back at least to the classical paper [12] by Eppstein, Paterson, and Yao in which, among many other results, they made the following simple observation: two edges with the same endpoint meet at an angle of at least $\pi/3$. Hence, the maximum indegree of the Nearest Neighbor Graph of $P\subset\mathbb{R}^{d}$ is bounded from above by the kissing number of $\mathbb{R}^{d}$ regardless of the size of $P$, see [7].
@@
For $n$ points on the line we can obtain the maximum indegree about $\log{n}$, which is optimal. (Unless specified otherwise, all logarithms in this paper are in base $2$.)
Theorem 1.1
@@
A similar maximum indegree (with a loss factor $1/(4d)$) can be obtained for $n$ points in $\mathbb{R}^{d}$, and in view of Theorem 1.1, this bound is best possible apart from the loss factor.
Theorem 1.2
@@
The maximum indegree we can obtain in a metric space on $n$ points is only about $\sqrt{\log{n}}$, and closing the gap between this lower bound and the $\lceil\log{n}\rceil$ upper bound remains as a challenging open problem.
Theorem 1.3
For every $n$-element metric space, there exists an order such that the corresponding ordered Nearest Neighbor Graph has maximum indegree
$\Omega\big{(}\sqrt{\log(n)/\log\log(n)}\big{)}$.
@@
Proof
Write $n=|P|$ and fix an arbitrary point $p_{n}\in P$. For $11\leq i\leq n-$ 1, let $p_{n-i}$ be the point closest to $p_{n-i+1}$ among the points $P\setminus\{p_{n},p_{n-1},\dots,p_{n-i+1}\}$. Now consider the order $\pi=(p_{1},p_{2},\dots,p_{n})$. It is easy to check that the ordered Nearest Neighbor Graph $G$ corresponding to $\pi$ is just the path $p_{n}\rightarrow p_{n-1}\rightarrow\ldots\rightarrow p_{1}$. Indeed, at the moment when $p_{i}$ is introduced, we connect $p_{i}$ to the point among $\{p_{1},p_{2},\dots,p_{i-1}\}$ that is closest to $p_{i}$, which is exactly $p_{i-1}$ according to our construction of $\pi$. ∎
2 Points on the line: proof of Theorem 1.1
Upper bound. We construct the desired set of points on the line recursively. For $n=2^{1}$, we put $P=P_{1}=\{0,1\}$. For $k\geq 1$ and $n=2^{k+1}$, we put $P=P_{k+1}=P_{k}\cup(3^{k}+P_{k})$. It is straightforward to verify by induction that ${\rm diam}(P_{k})=(3^{k}-1)/2$, which is smaller than the distance between the two ‘halves’ of $P_{k+1}$. Therefore, in every order of $P_{k+1}$, each point can have at most one edge incoming from the other ‘half’. Hence by induction, the indegree of each point does not exceed $k+1=\log{n}$. Finally, if $2^{k}<n\leq 2^{k+1}$, then an arbitrary subset $P\subset P_{k+1}$ of size $n$ inherits the latter property.
Lower bound. Choose $k$ such that $\lceil\log{n}\rceil=k+1$, i.e., $2^{k}<n\leq 2^{k+1}$. The proof is by induction on $k$. Note that the statement is trivial for $k=0$, so let us now assume that $k\geq 1$. We identify a suitable subset $P^{\prime}\subset P$ such that its reveal already forces the indegree of one of the points to reach the required threshold. The remaining points in $P\setminus P^{\prime}$ are introduced afterwards and clearly do not decrease any degree. We may assume that $P$ lies on the $x$-axis. We outline a recursive procedure ${\rm Order}$ for generating a suitable order that employs the following subroutines and a distinguished element of $P$:
@@ 3. Step 3: List $P$ as follows: ${\rm Center}(A)$, $b$, ${\rm Delete}({\rm Order}(A),{\rm Center}(A))$,
where the last inequality holds by induction hypothesis since $|A|\geq 2^{k-1}+1$. ∎
3 Points in $\mathbb{R}^{d}$: proof of Theorem 1.2
Here we further develop the ideas from the previous subsection to translate them from the line to $\mathbb{R}^{d}$, for any fixed $d$. As the main supplementary tool, we use the following standard result in discrete geometry, that can be deduced, e.g., from [19, Ineq. (4) and (6)].
Theorem 3.1
Let $K$ be a convex set in $\mathbb{R}^{d}$, $K_{0}$ be its interior, and $r<1$ be a positive number. Then $K$ can be covered by at most $(1+1/r)^{d}(d\ln d+d\ln\ln d+5d)$ translates of $-rK_{0}$, reflected scaled copies of $K_{0}$.
For a finite point set $P\subset\mathbb{R}^{d}$, we use ${\rm diam}(P)$ to denote its diameter, i.e. the maximum distance between any pair in $P$, and use ${\rm conv}(P)$ to denote its convex hull.
Corollary 1
Let $P$ be a finite set of points in $\mathbb{R}^{d}$ such that ${\rm diam}(P)\leq 1$. Then $P$ can be partitioned into at most $216^{d}/$ 2 subsets of diameter less than $1/2$.
Proof
We apply Theorem 3.1 with $K={\rm conv}(P)$ and $r=1/2$. Note that ${\rm diam}(K)={\rm diam}(P)\leq 1$, and thus the distance between any two points of $-(1/2)K_{0}$ is strictly less than $1/2$. It remains only to note that the number of translates given by Theorem 3.1 is smaller than $216^{d}/$ 2 for all $d\geq 2$. ∎
@@
  3. Step 3: By Corollary 1, partition $A$ into at most $216^{d}/$ 2 subsets where each subset has diameter less than $1/2$. One of these subsets $C\subset A$ contains at least $n/16^{d}$ points.
@@ 5. Step 5: ${\rm Center}(P)\leftarrow{\rm Center}(C)$
$\Delta^{-}(P)\geq\Delta^{-}(C)+1\geq\log_{16^{d}}{(n/16^{d})}+1=\log{n}/(4d)-1+1=\log{n}/(4d).\;\squareforqed$
4 Abstract metric spaces: proof of Theorem 1.3
Let $\rho$ be a metric on a finite set $P=\{p_{1},\dots,p_{n}\}$, namely a positive and symmetric function $\rho:P\times P\to\mathbb{R}_{+}$ satisfying the triangle inequality. So far we labeled the points of $P$ arbitrarily just for definiteness, the desired order on them will be constructed later. We shall assume the notion of hypergraph, and the unfamiliar readers are referred to, say [14], for related definitions.
We define a $3$-coloring of a complete $3$-uniform hypergraph $K_{n}^{(3)}$ on the vertex set $[n]$ as follows: for all $1\leq i_{1}<i_{2}<i_{3}\leq n$, we color the triple $\{i_{1},i_{2},i_{3}\}$ by red (resp. green or blue) whenever $v_{i_{2}}v_{i_{3}}$ (resp. $v_{i_{1}}v_{i_{3}}$ or $v_{i_{1}}v_{i_{2}}$) is the ‘shortest side’ of a triangle $v_{i_{1}}v_{i_{2}}v_{i_{3}}$, that is $\rho(v_{i_{2}},v_{i_{3}})<\rho(v_{i_{1}},v_{i_{2}})$ and $\rho(v_{i_{2}},v_{i_{3}})<\rho(v_{i_{1}},v_{i_{3}})$. Recall that our points are in general position, meaning that no point triple determines an isosceles triangle, and so each triple gets exactly one color. A subhypergraph of $K_{n}^{(3)}$ is monochromatic if all its triples are of the same color. A forward star $S_{k}^{(3)}$ is a 3-uniform hypergraph on an ordered vertex set $\{i_{1},\ldots,i_{k}\}\subset[n]$, labeled in the ascending order, consisting of edges $\{i_{1},i_{j},i_{j^{\prime}}\}$ for all $i_{j}<i_{j^{\prime}}$. We claim that if there exists a red clique $K_{k}^{(3)}$, or a green forward star $S_{k}^{(3)}$, or a blue forward star $S_{k}^{(3)}$, then there exists an order such that the corresponding Nearest Neighbor Graph has maximum indegree at least $k-1$. To verify this claim, suppose that such a structure exists on the vertex set $I=\{i_{1},\ldots,i_{k}\}$, labeled in the ascending order, and consider the following three cases.
@@
$v_{i_{k}},v_{i_{1}},v_{i_{2}},\dots,v_{i_{k-1}},{\rm AnyOrder}(P\setminus I),$
the indegree of $v_{i_{k}}$ is at least $k-1$. Indeed, note that as each new vertex $v_{i_{j}}$ is revealed, $\rho(v_{i_{j}},v_{i_{k}})$ is the shortest distance among all currently visible points. Hence, $v_{i_{j}}\to v_{i_{k}}$ is an edge of the ordered Nearest Neighbor Graph for all $11\leq j\leq k-$ 1.
@@
$v_{i_{1}},v_{i_{k}},v_{i_{k-1}},\dots,v_{i_{2}},{\rm AnyOrder}(P\setminus I),$
is at least $k-1$. Indeed, when $v_{i_{j}}$ is revealed, the distance $\rho(v_{i_{1}},v_{i_{j}})$ will be the shortest among all current points as $\{i_{1},i_{j},i_{\ell}\}$ is always blue.
@@
$v_{i_{1}},v_{i_{2}},v_{i_{3}},\dots,v_{i_{k}},{\rm AnyOrder}(P\setminus I),$
is at least $k-1$. Indeed, when $v_{i_{j}}$ is revealed, an arbitrary currently visible point $v_{i_{\ell}}$ satisfies $\{i_{1},i_{\ell},i_{j}\}$ being green. So the distance $\rho(v_{i_{1}},v_{i_{j}})$ will be the shortest.
The last ingredient of the proof is the following Ramsey type result due to He and Fox; more precisely, their Theorem 1.2 and Lemma 5.2 in  [14]. We rewrite this proof to make it explicit that the argument in [14] produces forward stars and also extends to more than two colors.
Theorem 4.1
@@
Applying this theorem with $k=c^{\prime}\sqrt{\log(n)/\log\log(n)}$ for a sufficiently small $c^{\prime}>0$, we ensure the existence of a monochromatic special structure, and thus the existence of an order such that the corresponding ordered Nearest Neighbor Graph has maximum indegree at least $k-1$, which completes the proof of Theorem 1.3.
Proof (of Theorem 4.1)
@@
Now we describe how we create edges: once a new vertex $v$ is picked, we keep creating edges from $v$ to the red vertices in $U$ following the order until we encounter a green or blue edge; if the last edge we created is green (resp. blue), we color the vertex green (resp. blue); otherwise we color this vertex red. The new vertex and the created edges are considered as added into $G$. The next iteration begins as long as there are still vertices left in $W$.
Notice that if there are $k$ red vertices in $G$, then by construction all the 2-edges between them are red. And again by the way we color the edges in $G$, these vertices correspond to a 3-uniform red clique $K_{k}^{(3)}$ inside $K_{n}^{(3)}$. Hence, without loss of generality, we assume that there are fewer than $k$ red vertices in $G$. Given this assumption, if there are $(k-1)^{2}$ green vertices in $G$, then by our construction and the pigeonhole principle, there must be one red vertex $v_{1}$ and $k-1$ green vertices $v_{2},\dots,v_{k}$ such that $\{v_{1},v_{i}\}$ is green for $2\leq i\leq k$. Then by the way we color the edges in $G$, these vertices correspond to a 3-uniform green forward star $S_{k}^{(3)}$ inside $K_{n}^{(3)}$. Hence, without loss of generality, there are fewer than $(k-1)^{2}$ green vertices in $G$, and a similar claim holds for blue vertices as well. Therefore, $G$ has fewer than $k+2(k-1)^{2}$ vertices in total.
Next, we upper bound the number of edges in $G$. Since each green vertex gives exactly one green edge and there is no other source of green edges, there are fewer than $(k-1)^{2}$ green edges. Similarly, there are fewer than $(k-1)^{2}$ blue edges. There are at most $(k-1)^{2}$ red edges between red vertices. To count other red edges, notice that the $i$-th red vertex has fewer than $k-1$ green (resp. blue) edges adjacent to it, and each endpoint of these edges is adjacent to $i-1$ red edges. In summary, the number of red edges in $G$ is fewer than
@@
Finally, notice that each vertex in $G$ decreases the size of $W$ by 1, each green or blue edge shrinks the size of $W$ by $1/k$, and each red edge shrinks the size of $W$ by $(1-2/k)$. Since by the end of our construction of $G$ when $W$ is depleted, $G$ has fewer than $k+2(k-1)^{2}$ vertices, fewer than $2(k-1)^{2}$ green or blue edges, and fewer than $k(k-1)^{2}/2$ red edges, the size of $W$ in the beginning, namely $n$, is at most

### v3

@@ 1 Introduction
For a given point set $P$ in the plane, in $d$-dimensional Euclidean space, or, even more generally, in an arbitrary metric space, its Nearest Neighbor Graph is a directed graph based on minimum distances. More precisely, these points are the vertices of the Nearest Neighbor Graph, and each vertex has precisely one outgoing edge towards its closest neighbor. To make this notion well-defined, we assume that our point set is in general position, which here means that no three points determine an isosceles triangle.
These graphs play an important role in modern Computational Geometry due to their relevance in the context of computing geometric shortest paths [16, 17]. Further applications of Nearest Neighbor Graphs for computing spanners, well-separated pairs, and approximate minimum spanning trees in $\mathbb{R}^{d}$ can be found in the survey [21], books [10, 19], or monograph [18], see also [14]. A systematic study of the basic combinatorial properties of Nearest Neighbor Graphs dates back at least to the classical paper [13] by Eppstein, Paterson, and Yao in which, among many other results, they made the following simple observation: two edges with the same endpoint meet at an angle of at least $\pi/3$. Hence, the maximum indegree of the Nearest Neighbor Graph of a point set $P\subset\mathbb{R}^{d}$ is bounded from above by the kissing number of $\mathbb{R}^{d}$, see [8], i.e., the maximum number of non-overlapping unit spheres that can touch a unit sphere in $\mathbb{R}^{d}$, regardless of the size of $P$.
@@
For $n$ points on the line, the maximum indegree can be made $\lceil\log n\rceil$, which is optimal. (Unless specified otherwise, all logarithms in this paper are in base $2$.)
Theorem 1.
@@
We obtain a similar maximum indegree (with a loss factor $1/(4d)$) for $n$ points in $\mathbb{R}^{d}$, and in view of Theorem 1, this bound is best possible apart from the loss factor.
Theorem 2.
@@
The maximum indegree we manage to obtain in a metric space on $n$ points is only about $\sqrt{\log{n}}$, and closing the gap between this lower bound and the $\lceil\log{n}\rceil$ upper bound remains as a challenging open problem.
Theorem 3.
For every $n$-element metric space, there exists an order such that the corresponding ordered Nearest Neighbor Graph has maximum indegree $\Omega\big(\sqrt{\log(n)/\log\log(n)}\big)$.
@@
Proof.
Write $n=|P|$ and fix an arbitrary point $p_{n}\in P$. For $1\leq i\leq n-1$, let $p_{n-i}$ be the point closest to $p_{n-i+1}$ among the points $P\setminus\{p_{n},p_{n-1},\dots,p_{n-i+1}\}$. Now consider the order $\pi=(p_{1},p_{2},\dots,p_{n})$. It is easy to check that the ordered Nearest Neighbor Graph $G$ corresponding to $\pi$ is just the path $p_{n}\rightarrow p_{n-1}\rightarrow\ldots\rightarrow p_{1}$. Indeed, at the moment when $p_{i}$ is introduced, we connect $p_{i}$ to the point among $\{p_{1},p_{2},\dots,p_{i-1}\}$ that is closest to $p_{i}$, which is exactly $p_{i-1}$ according to our construction of $\pi$. ∎
Notation.
For a finite point set $P\subset\mathbb{R}^{d}$, we use ${\rm diam}(P)$ to denote its diameter, i.e., the maximum distance between any pair of points in $P$, and ${\rm conv}(P)$ to denote its convex hull.
2 Points on the line: proof of Theorem 1
Upper bound. We construct the desired set of points on the line recursively. For $n=2^{1}$, we put $P=P_{1}=\{0,1\}$. For $k\geq 1$ and $n=2^{k+1}$, we put $P=P_{k+1}=P_{k}\cup(3^{k}+P_{k})$. It is straightforward to verify by induction that ${\rm diam}(P_{k})=(3^{k}-1)/2$, which is smaller than the distance between the two ‘halves’ of $P_{k+1}$, namely, $P_{k}$ and $3^{k}+P_{k}$. Indeed, ${\rm diam}(P_{k})=(3^{k}-1)/2<(3^{k}+1)/2$. Therefore, in every order of $P_{k+1}$, each point can have at most one edge incoming from the other ‘half’. Hence by induction, the indegree of each point does not exceed $k+1=\log{n}$. Finally, if $2^{k}<n\leq 2^{k+1}$, then an arbitrary subset $P\subseteq P_{k+1}$ of size $n$ inherits the latter property.
Lower bound. Choose $k$ such that $\lceil\log{n}\rceil=k+1$, i.e., $2^{k}<n\leq 2^{k+1}$. The proof is by induction on $k$. Note that the statement is trivial for $k=0$, so let us now assume that $k\geq 1$. We identify a suitable subset $P^{\prime}\subseteq P$ such that its reveal already forces the indegree of one of the points to reach the required threshold. The remaining points in $P\setminus P^{\prime}$ are introduced afterwards and clearly do not decrease any degree. We may assume that $P$ lies on the $x$-axis. Let $X\subseteq P$. We outline a recursive procedure $\rm Order$} for generating a suitable order that employs the following subroutines and a distinguished element of $P$:
@@ 3. Step 3: List $P$ as follows: ${\rm Center}(A)$, $b$, ${\rm Delete}({\rm Order}(A),{\rm Center}(A))$,
where the last inequality holds by induction hypothesis since $2^{k-1}<|A|$. ∎
3 Points in $\mathbb{R}^{d}$: proof of Theorem 2
Here we further develop the ideas from the previous subsection to translate them from the line to $\mathbb{R}^{d}$, for any fixed $d$. As the main supplementary tool, we use the following standard result in discrete geometry, that can be deduced, e.g., from [20, Ineq. (4) and (6)], see also [4]
Theorem 4.
Let $K$ be a convex set in $\mathbb{R}^{d}$, $K_{0}$ be its interior, and $r<1$ be a positive number. Then $K$ can be covered by at most $(1+1/r)^{d}(d\ln d+d\ln\ln d+5d)$ translates of $rK_{0}$, reflected scaled copies of $K_{0}$.
Corollary 1.
Let $P$ be a finite set of points in $\mathbb{R}^{d}$ such that ${\rm diam}(P)\leq 1$. Then $P$ can be partitioned into at most $16^{d}/2$ subsets of diameter less than $1/2$.
Proof.
We apply Theorem 4 with $K={\rm conv}(P)$ and $r=1/2$. Note that ${\rm diam}(K)={\rm diam}(P)\leq 1$, and thus the distance between any two points of $(1/2)K_{0}$ is strictly less than $1/2$. It remains only to note that the number of translates given by Theorem 4 is smaller than $16^{d}/2$ for all $d\geq 2$. ∎
@@ 3 Points in $\mathbb{R}^{d}$: proof of Theorem 2
  3. Step 3: By Corollary 1, partition $A$ into at most $16^{d}/2$ subsets where each subset has diameter less than $1/2$. One of these subsets $C\subseteq A$ contains at least $n/16^{d}$ points.
@@ 5. Step 5: ${\rm Center}(P)\leftarrow{\rm Center}(C)$
$\Delta^{-}(P)\geq\Delta^{-}(C)+1\geq\log_{16^{d}}{(n/16^{d})}+1=\log{n}/(4d)-1+1=\log{n}/(4d).$$\qed$
4 Abstract metric spaces: proof of Theorem 3
Let $\rho$ be a metric on a finite set $V=\{v_{1},\dots,v_{n}\}$, namely a positive and symmetric function $\rho:V\times V\to\mathbb{R}_{+}$ satisfying the triangle inequality. So far we labeled the points of $V$ arbitrarily just for definiteness, the desired order on them will be constructed later. We assume familiarity with hypergraphs, otherwise readers are referred to [15] for related definitions.
We define a $3$-coloring of a complete $3$-uniform hypergraph $K_{n}^{(3)}$ on the vertex set $[n]$ as follows: for all $1\leq i_{1}<i_{2}<i_{3}\leq n$, we color the triple $\{i_{1},i_{2},i_{3}\}$ by red (resp. green or blue) whenever $v_{i_{2}}v_{i_{3}}$ (resp. $v_{i_{1}}v_{i_{3}}$ or $v_{i_{1}}v_{i_{2}}$) is the ‘shortest side’ of a triangle $v_{i_{1}}v_{i_{2}}v_{i_{3}}$, that is $\rho(v_{i_{2}},v_{i_{3}})<\rho(v_{i_{1}},v_{i_{2}})$ and $\rho(v_{i_{2}},v_{i_{3}})<\rho(v_{i_{1}},v_{i_{3}})$. Recall that our points are in general position, meaning that no point triple determines an isosceles triangle, and so each triple gets exactly one color. A subhypergraph of $K_{n}^{(3)}$ is monochromatic if all its triples are of the same color. A forward star $S_{k}^{(3)}$ is a 3-uniform hypergraph on an ordered vertex set $\{i_{1},\ldots,i_{k}\}\subseteq[n]$, labeled in the ascending order, consisting of edges $\{i_{1},i_{j},i_{j^{\prime}}\}$ for all $i_{j}<i_{j^{\prime}}$. We claim that if there exists a red clique $K_{k}^{(3)}$, or a green forward star $S_{k}^{(3)}$, or a blue forward star $S_{k}^{(3)}$, then there exists an order such that the corresponding Nearest Neighbor Graph has maximum indegree at least k1k-1. To verify this claim, suppose that such a structure exists on the vertex set $I=\{i_{1},\ldots,i_{k}\}$, labeled in the ascending order, and consider the following three cases.
@@ 4 Abstract metric spaces: proof of Theorem 3
$v_{i_{k}},v_{i_{1}},v_{i_{2}},\dots,v_{i_{k-1}},{\rm AnyOrder}(V\setminus I),$
the indegree of $v_{i_{k}}$ is at least k1k-1. Indeed, note that as each new vertex $v_{i_{j}}$ is revealed, $\rho(v_{i_{j}},v_{i_{k}})$ is the shortest distance among all currently visible points. Hence, $v_{i_{j}}\to v_{i_{k}}$ is an edge of the ordered Nearest Neighbor Graph for all $1\leq j\leq k-1$.
@@
$v_{i_{1}},v_{i_{k}},v_{i_{k-1}},\dots,v_{i_{2}},{\rm AnyOrder}(V\setminus I),$
is at least k1k-1. Indeed, when $v_{i_{j}}$ is revealed, the distance $\rho(v_{i_{1}},v_{i_{j}})$ will be the shortest among all current points as $\{i_{1},i_{j},i_{\ell}\}$ is always blue.
@@
$v_{i_{1}},v_{i_{2}},v_{i_{3}},\dots,v_{i_{k}},{\rm AnyOrder}(V\setminus I),$
is at least k1k-1. Indeed, when $v_{i_{j}}$ is revealed, an arbitrary currently visible point $v_{i_{\ell}}$ satisfies $\{i_{1},i_{\ell},i_{j}\}$ being green. So the distance $\rho(v_{i_{1}},v_{i_{j}})$ will be the shortest.
The last ingredient of the proof is the following Ramsey-type result due to He and Fox; more precisely, their Theorem 1.2 and Lemma 5.2 in  [15]. We rewrite this proof to make it explicit that the argument in [15] produces forward stars and also extends to more than two colors.
Theorem 5.
@@
Applying this theorem with $k=c^{\prime}\sqrt{\log(n)/\log\log(n)}$ for a sufficiently small $c^{\prime}>0$, we ensure the existence of a monochromatic special structure, and thus the existence of an order such that the corresponding ordered Nearest Neighbor Graph has maximum indegree at least k1k-1, which completes the proof of Theorem 3.
Proof of Theorem 5.
@@
Now we describe how we create edges: once a new vertex $v$ is picked, we keep creating edges from $v$ to the red vertices in $U$ following the order until we encounter a green or blue edge; if the last edge we created is green (resp. blue), we color the vertex green (resp. blue); otherwise we color this vertex red. The new vertex and the created edges are considered as added into $G$. By construction, every edge in $G$ leads to a red vertex and each green (resp. blue) vertex is adjacent to exactly one green (resp. blue) edge. The next iteration begins as long as there are still vertices left in $W$.
If there are $k$ red vertices in $G$, then by construction all the 2-edges between them are red. And again by the way we color the edges in $G$, these vertices correspond to a 3-uniform red clique $K_{k}^{(3)}$ inside $K_{n}^{(3)}$. Hence, without loss of generality, we may assume that there are fewer than $k$ red vertices in $G$. Given this assumption, if there are $(k-1)^{2}$ green vertices in $G$, then by our construction and the pigeonhole principle, there must be one red vertex $v_{1}$ and k1k-1 green vertices $v_{2},\dots,v_{k}$ such that $\{v_{1},v_{i}\}$ is green for $2\leq i\leq k$. Then by the way we color the edges in $G$, these vertices correspond to a 3-uniform green forward star $S_{k}^{(3)}$ inside $K_{n}^{(3)}$. Hence, without loss of generality, there are fewer than $(k-1)^{2}$ green vertices in $G$, and a similar claim holds for blue vertices as well. Therefore, $G$ has fewer than $k+2(k-1)^{2}$ vertices in total.
Next, we upper bound the number of edges in $G$. Since green edges only originate from green vertices, there are fewer than $(k-1)^{2}$ green edges. Similarly, there are fewer than $(k-1)^{2}$ blue edges. There are at most $(k-1)^{2}$ red edges between red vertices. To count other red edges, notice that the $i$-th red vertex has fewer than k1k-1 green (resp. blue) edges adjacent to it, and each endpoint of these edges is adjacent to i1i-1 red edges. In summary, the number of red edges in $G$ is fewer than
@@
Finally, notice that each vertex in $G$ decreases the size of $W$ by 1, each green or blue edge preserves at least a $1/k$-fraction of the elements of $W$ and each red edge preserves at least a (12/k)(1-2/k)-fraction of the elements of $W$. Since by the end of our construction of $G$ when $W$ is depleted, $G$ has fewer than $k+2(k-1)^{2}$ vertices, fewer than $2(k-1)^{2}$ green or blue edges, and fewer than $k(k-1)^{2}/2$ red edges, the size of $W$ in the beginning, namely $n$, is at most